# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 9.6. Splitting stars and torsos
Assume not for a contradiction, and let $v$ be a neighbour of $u$ in the $D$-link. Recall that $u\notin D$ by Lemma 9.42. So the edge $uv$ dangles from the $D$-link through the $A$-link. Hence the corner $AD$ is potter by the (Crossing Lemma). (4.1), and the triangle-torso $H_{o(A)}$ has $uv$ as free edge. By assumption, $o(A)$ and $s$ are not adjacent in $O$. By Lemma 9.15, there is precisely one torso $H_{r}$ in between the two triangle-torsos $H_{o(A)}$ and $H_{s}$, which also contains $u$. But $H_{r}$ has $u$ as tip since $uv$ is free in $H_{o(A)}$, which by Lemma 9.15 contradicts that $H_{s}$ also has $u$ as tip. ∎
Figure 9.8. The situation in Lemma 9.44.
Let $P$ be a path, and let $u,v$ be vertices on $P$. Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$. The statement of the following lemma is supported by Figure 9.8.
@@
Assume the (Tutte-bagel setting). (9.12). Let $P$ and $Q$ be the two paths that are the two components of $O-o(C)-o(D)$. Assume that $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $s$ be a node in the same component of $P-o(A)$ as $c$ such that $H_{s}$ is a 4-cycle or a triangle. Assume that if the $A$-link is represented by a triangle-torso, then $s$ is not adjacent to $o(A)$ in $P$. Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$. Then:
@@
    $S(A^{\prime},B^{\prime})=S\cup(D$-link)))), where $S$ consists of the two free edges of $H_{s}$ if $H_{s}$ is a 4-cycle and of the free edge and the tip of $H_{s}$ if $H_{s}$ is a triangle.
@@
(ii) holds by construction, and gives that the mixed-separation $(A^{\prime},B^{\prime})$ has order four. We show (i) next.
Matching-condition. Let $ab$ be an edge in the $D$-link with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Let $f$ be an edge in $S$. Assume for a contradiction that $ab$ and $f$ share an endvertex, which can only be $a$. Since $a\in(A\smallsetminus B)\cap D$ and $V(G_{s})\subseteq A\cap C$, we find that $a$ is in the $A$-link. So $ab$ dangles from the $A$-link through the $D$-link. Then the corner $AD$ is potter by the (Crossing Lemma). (4.1). Thus the $A$-link is represented by a triangle-torso, so $a\notin D$ by Lemma 9.42, a contradiction.
Degree-condition. It is straightforward to check that every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $B^{\prime}\smallsetminus A^{\prime}$. Now let $v\in S(A^{\prime},B^{\prime})$ be a vertex; we claim that it has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$. If $v$ is in $S$, then $H_{s}$ is a triangle with tip $v$. Let $H_{r}$ denote the torso neighbouring $H_{s}$ in the direction of $o(A)$ on the path $P$. If $H_{r}$ is 3-connected, then $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $G_{s}\subseteq G[A\cap C]$. Otherwise $H_{r}$ is a triangle, and the adhesion-graph in between $H_{r}$ and $H_{s}$ is a $K_{2}$-subgraph of $G$ by Lemma 9.4. Hence $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $r\neq o(A)$ by assumption. So we may assume next that $v$ lies in the $D$-link. Every neighbour of $v$ in $A\smallsetminus B$ also lies in $A^{\prime}\smallsetminus B^{\prime}$, except possibly when a neighbour of $v$ lies in $S$ or an edge incident to $v$ lies in $S$. The former exception is impossible by Lemma 9.43, while the latter is excluded by Lemma 9.42. This conclude the proof of (i).
(iii). We have to find two vertices in $B\cup D$ that are contained in $A^{\prime}\smallsetminus B^{\prime}$ and in $B^{\prime}\smallsetminus A^{\prime}$, respectively. Every vertex in $B\smallsetminus A$ lies in the intersection of $B\cup D$ with $B^{\prime}\smallsetminus A^{\prime}$, and $B\smallsetminus A$ is non-empty. Any vertex in $D\smallsetminus B$ lies in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma 9.42, so it remains to show that $D\smallsetminus B$ is non-empty. If $D\smallsetminus B$ is empty, then the $D$-link consists of two edges. But both these edges have ends in the corner for $AD$ by the (Crossing Lemma). (4.1), a contradiction. ∎
@@
Assume the (Tutte-bagel setting). (9.12). Let $P$ and $Q$ be the two paths that are the two components of $O-o(C)-o(D)$. Assume that $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $e$ be an edge of $O$ such that the torsos at the ends of $e$ are 3-connected. Assume that $e$ stems from the $AC$-Tutte-path. Let $P^{\prime}:=dP\mathring{e}$ and $Q^{\prime}:=\mathring{e}Pc\cup Q$. Then:
@@
    $S(A^{\prime},B^{\prime})=\pi\cup(D$-link)))), where $\pi$ is the adhesion-set of $ac$.
@@
The proof is analogous to the proof of Lemma 9.44, except that everything is straightforward because the two 3-connected torsos at the ends of $e$ both have disjoint adhesion-sets and have all their vertices contained in $A\cap C$. ∎
@@
Assume the (Tutte-bagel setting). (9.12). Let $i:=1$ if $H_{t(1)}$ is 3-connected, and $i:=2$ otherwise. Similarly, let $j:=-1$ if $H_{t(-1)}$ is 3-connected, and $j:=-2$ otherwise. The right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$ has sides
@@
This is just a different perspective on Lemma 9.14. ∎
@@
Assume the (Tutte-bagel setting). (9.12). Let $(E,F)$ be a totally-nested tetra-separation of $G$ that interlaces $\sigma$. Then there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ or $(F,E)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
@@
Let us show that $x$ is not the only vertex of $T^{\prime}$. Otherwise the torso $H_{x}$ is a triangle or 4-cycle. By Lemma 9.18, $H_{x}$ is a triangle and both the $A$-link and the $C$-link are represented by triangle-torsos. In particular, $A\cap C=3$. But $E$, as a side of a tetra-separation, has size $\geqslant 4$, and $E$ is included in $A\cap C$, a contradiction. So $x$ is not the only vertex of $T^{\prime}$.
Let $y$ be an endvertex of $T^{\prime}$ that is distinct from $x$. By symmetry, we may assume that $y$ is adjacent or incident to $o(A)$ when viewed as a node of $O$. Then the (Link-Moving Lemma for vertices). (9.44) produces a $G$-lift. This concludes the proof that 7 holds.
For every admissible $(M,N)$ we use 7 to choose a $G$-lift $(\bar{M},\bar{N})$ so that for $(N,M)$ we choose $(\bar{N},\bar{M})$. We claim that exactly one of $E\subseteq\bar{M}$ and $E\subseteq\bar{N}$ holds, for all $(M,N)$. Indeed, otherwise $(\bar{M},\bar{N})$ separates $E$, and since it also separates $F$ by definition, $(\bar{M},\bar{N})$ crosses $(E,F)$, a contradiction. We let
$\omega$ :=${}:=\{\,(M,N):(M,N)\text{ is admissable and }\bar{N}\supseteq E\,\}$
$\Omega$ :=${}:=\bigcap\,\{\,N\smallsetminus M:(M,N)\in\omega\,\}\cap V(T^{\prime}).$
@@
Suppose first that $\omega$ is empty. Then $V(T^{\prime})$ is empty. Hence $T$ is a $K_{2}$ both whose endvertices have triangle or 4-cycle torsos that represent links. By Lemma 9.4, both torsos are triangles. Hence the corner $AC$ is potter, so $A\cap C=2$, contradicting that $E\subseteq A\cap C$ where $E$ has size $\geqslant 4$.
@@
To show that $(E,F)\leqslant$ the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$, by Lemma 9.46 and Lemma 5.7 it remains to show $E\subseteq V(G_{t^{*}})$. Assume for a contradiction that there is a vertex $v\in E\smallsetminus V(G_{t^{*}})$. Recall that $E\subseteq A\cap C$, so in particular $E\subseteq V(\beta)$. Hence there is a node $s$ of $T$ with $v\in V_{s}$, and $s\neq t^{*}$ by assumption. Let $r$ be the penultimate node on the path $sTt^{*}$.
Assume first that the torso $H_{r}$ is 3-connected. Then the edge $rt^{*}$ is admissible. Let $(M,N)\in\omega$ with $S(M,N)=\{rt^{*}\}$. Note that $r\in M\smallsetminus N$ and $t^{*}\in N\smallsetminus M$ since $\{t^{*}\}=\Omega$. Let $(\hat{M},\hat{N})$ be the $\beta$-lift of $(M,N)$. Observe that $v\in\hat{M}\smallsetminus\hat{N}$. Therefore, $v\in\bar{M}\smallsetminus\bar{N}$, contradicting that $E\subseteq\bar{N}$.
Assume otherwise that the torso $H_{r}$ is either a triangle or a 4-cycle. Then $r$ is admissible or $r$ is one of $o(A)$ and $o(C)$. Assume first that $r$ is admissible. Let $(M,N)\in\omega$ with $S(M,N)=\{r\}$. Note that $r,s\in M$ and $t^{*}\in N\smallsetminus M$. Let $(\hat{M},\hat{N})$ be the $\beta$-lift of $(M,N)$. Then $v\in\hat{M}\smallsetminus\hat{N}$, so $v\in\bar{M}\smallsetminus\bar{N}$, contradicting that $E\subseteq\bar{N}$. Assume otherwise that $r=o(A)$, say. Then $H_{r}$ is a triangle or 4-cycle representing the $A$-link. Since $v$ does not lie in $H_{t^{*}}$, we in particular have that $v$ does not lie in the adhesion-set of the edge $rt^{*}$. But then $v$ lies in $D\smallsetminus C$, contradicting that $E\subseteq C$. ∎
@@
Assume for a contradiction that $\sigma$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$. By Lemma 9.48, there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, say, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
Assume first that $H_{t}$ is good. Then $(U_{t},W_{t})$ is a totally-nested tetra-separation by 9.41. In particular, $(U_{t},W_{t})\in\sigma$. Since $(E,F)$ interlaces $\sigma$, we have one of $(U_{t},W_{t})<(E,F)$ and $(U_{t},W_{t})<(F,E)$. We cannot have $(U_{t},W_{t})<(F,E)$, since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t})$ contradicts the fact that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by 5.2. Hence $(U_{t},W_{t})<(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$. Since $(U_{t},W_{t})$ is the left-shift of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
Assume otherwise that $G_{t}$ is bad. Hence $G_{t}$ has four or five vertices. If $H_{t}$ has four vertices, then all but at most one of the four vertices in $S(U^{\prime}_{t},W^{\prime}_{t})$ has $\geqslant 2$ neighbours in $W^{\prime}_{t}\smallsetminus U^{\prime}_{t}$. If $H_{t}$ has five vertices, then all four vertices in $S(U^{\prime}_{t},W^{\prime}_{t})$ have $\geqslant 2$ neighbours in $W^{\prime}_{t}\smallsetminus U^{\prime}_{t}$. In either case, we have $|U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}|\leqslant 1$. Since $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, we have $E\smallsetminus F\subseteq U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$. Therefore, $|E\smallsetminus F|\leqslant 1$, which by Lemma 5.5 contradicts that $(E,F)$ is a tetra-separation. ∎
@@
By the (Crossing Lemma). (4.1), all links have size two. By Lemma 9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$. By 9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. By Lemma 9.49, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$. Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$. Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$. The torso $\tau$ is 4-connected by Lemma 8.36. We can obtain $\tau$ from $\mathcal{O}$ by replacing each good torso with a $K_{4}$, leaving only 3-connected torsos with five or four vertices, and contracting some free edges in torsos that are 4-cycles or triangles. Each torso that is a 4-cycle neighbours two 3-connected torsos by Lemma 9.4, and the 4-cycle-torso witnesses that both 3-connected torsos are good. Hence the free edges of the 4-cycle-torsos are contained in the separators of the elements of $\sigma(\mathcal{O})$ that stem from the good neighbouring torsos, and so they are contracted. Hence $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. ∎

### v2

@@ 9.6. Splitting stars and torsos
Assume not for a contradiction, and let $v$ be a neighbour of $u$ in the $D$-link. Recall that $u\notin D$ by Lemma˜9.42. So the edge $uv$ dangles from the $D$-link through the $A$-link. Hence the corner $AD$ is potter by the (Crossing Lemma). (4.1), and the triangle-torso $H_{o(A)}$ has $uv$ as free edge. By assumption, $o(A)$ and $s$ are not adjacent in $O$. By Lemma˜9.15, there is precisely one torso $H_{r}$ in between the two triangle-torsos $H_{o(A)}$ and $H_{s}$, which also contains $u$. But $H_{r}$ has $u$ as tip since $uv$ is free in $H_{o(A)}$, which by Lemma˜9.15 contradicts that $H_{s}$ also has $u$ as tip. ∎
Figure 9.8. The situation in Lemma˜9.44.
Let $P$ be a path, and let $u,v$ be vertices on $P$. Recall that $uP\mathring{v}$ denotes the subpath of $P$ that starts at $u$ and ends just before the vertex $v$. The statement of the following lemma is supported by Figure˜9.8.
@@
Assume the (Tutte-bagel setting). (9.12). Let $P$ and $Q$ be the two paths that are the two components of Oo(C)o(D)O-o(C)-o(D). Assume that $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $s$ be a node in the same component of Po(A)P-o(A) as $c$ such that $H_{s}$ is a 4-cycle or a triangle. Assume that if the $A$-link is represented by a triangle-torso, then $s$ is not adjacent to $o(A)$ in $P$. Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$. Then:
@@
    $S(A^{\prime},B^{\prime})=S\cup(D$-link)), where $S$ consists of the two free edges of $H_{s}$ if $H_{s}$ is a 4-cycle and of the free edge and the tip of $H_{s}$ if $H_{s}$ is a triangle.
@@
˜(ii) holds by construction, and gives that the mixed-separation $(A^{\prime},B^{\prime})$ has order four. We show ˜(i) next.
Matching-condition. Let $ab$ be an edge in the $D$-link with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. Let $f$ be an edge in $S$. Assume for a contradiction that $ab$ and $f$ share an endvertex, which can only be $a$. Since $a\in(A\smallsetminus B)\cap D$ and $V(G_{s})\subseteq A\cap C$, we find that $a$ is in the $A$-link. So $ab$ dangles from the $A$-link through the $D$-link. Then the corner $AD$ is potter by the (Crossing Lemma). (4.1). Thus the $A$-link is represented by a triangle-torso, so $a\notin D$ by Lemma˜9.42, a contradiction.
Degree-condition. It is straightforward to check that every vertex in $S(A^{\prime},B^{\prime})$ has $\geqslant 2$ neighbours in $B^{\prime}\smallsetminus A^{\prime}$. Now let $v\in S(A^{\prime},B^{\prime})$ be a vertex; we claim that it has $\geqslant 2$ neighbours in $A^{\prime}\smallsetminus B^{\prime}$. If $v$ is in $S$, then $H_{s}$ is a triangle with tip $v$. Let $H_{r}$ denote the torso neighbouring $H_{s}$ in the direction of $o(A)$ on the path $P$. If $H_{r}$ is 3-connected, then $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $G_{s}\subseteq G[A\cap C]$. Otherwise $H_{r}$ is a triangle, and the adhesion-graph in between $H_{r}$ and $H_{s}$ is a $K_{2}$-subgraph of $G$ by Lemma˜9.4. Hence $v$ has two neighbours in $G_{r}$, both of which lie in $A^{\prime}\smallsetminus B^{\prime}$ since $r\neq o(A)$ by assumption. So we may assume next that $v$ lies in the $D$-link. Every neighbour of $v$ in $A\smallsetminus B$ also lies in $A^{\prime}\smallsetminus B^{\prime}$, except possibly when a neighbour of $v$ lies in $S$ or an edge incident to $v$ lies in $S$. The former exception is impossible by Lemma˜9.43, while the latter is excluded by Lemma˜9.42. This conclude the proof of ˜(i).
˜(iii). We have to find two vertices in $B\cup D$ that are contained in $A^{\prime}\smallsetminus B^{\prime}$ and in $B^{\prime}\smallsetminus A^{\prime}$, respectively. Every vertex in $B\smallsetminus A$ lies in the intersection of $B\cup D$ with $B^{\prime}\smallsetminus A^{\prime}$, and $B\smallsetminus A$ is non-empty. Any vertex in $D\smallsetminus B$ lies in $A^{\prime}\smallsetminus B^{\prime}$ by Lemma˜9.42, so it remains to show that $D\smallsetminus B$ is non-empty. If $D\smallsetminus B$ is empty, then the $D$-link consists of two edges. But both these edges have ends in the corner for $AD$ by the (Crossing Lemma). (4.1), a contradiction. ∎
@@
Assume the (Tutte-bagel setting). (9.12). Let $P$ and $Q$ be the two paths that are the two components of Oo(C)o(D)O-o(C)-o(D). Assume that $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $e$ be an edge of $O$ such that the torsos at the ends of $e$ are 3-connected. Assume that $e$ stems from the $AC$-Tutte-path. Let $P^{\prime}:=dP\mathring{e}$ and $Q^{\prime}:=\mathring{e}Pc\cup Q$. Then:
@@
    $S(A^{\prime},B^{\prime})=\pi\cup(D$-link)), where $\pi$ is the adhesion-set of $ac$.
@@
The proof is analogous to the proof of Lemma˜9.44, except that everything is straightforward because the two 3-connected torsos at the ends of $e$ both have disjoint adhesion-sets and have all their vertices contained in $A\cap C$. ∎
@@
Assume the (Tutte-bagel setting). (9.12). Let $i:=1$ if $H_{t(1)}$ is 3-connected, and $i:=2$ otherwise. Similarly, let j:=1j:=-1 if $H_{t(-1)}$ is 3-connected, and j:=2j:=-2 otherwise. The right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$ has sides
@@
This is just a different perspective on Lemma˜9.14. ∎
@@
Assume the (Tutte-bagel setting). (9.12). Let $(E,F)$ be a totally-nested tetra-separation of $G$ that interlaces $\sigma$. Then there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ or $(F,E)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
@@
Let us show that $x$ is not the only vertex of $T^{\prime}$. Otherwise the torso $H_{x}$ is a triangle or 4-cycle. By Lemma˜9.18, $H_{x}$ is a triangle and both the $A$-link and the $C$-link are represented by triangle-torsos. In particular, $A\cap C=3$. But $E$, as a side of a tetra-separation, has size $\geqslant 4$, and $E$ is included in $A\cap C$, a contradiction. So $x$ is not the only vertex of $T^{\prime}$.
Let $y$ be an endvertex of $T^{\prime}$ that is distinct from $x$. By symmetry, we may assume that $y$ is adjacent or incident to $o(A)$ when viewed as a node of $O$. Then the (Link-Moving Lemma for vertices). (9.44) produces a $G$-lift. This concludes the proof that ˜7 holds.
For every admissible $(M,N)$ we use ˜7 to choose a $G$-lift $(\bar{M},\bar{N})$ so that for $(N,M)$ we choose $(\bar{N},\bar{M})$. We claim that exactly one of $E\subseteq\bar{M}$ and $E\subseteq\bar{N}$ holds, for all $(M,N)$. Indeed, otherwise $(\bar{M},\bar{N})$ separates $E$, and since it also separates $F$ by definition, $(\bar{M},\bar{N})$ crosses $(E,F)$, a contradiction. We let
$\omega$ ${}:=\{\,(M,N):(M,N)\text{ is admissable and }\bar{N}\supseteq E\,\}$
$\Omega$ ${}:=\bigcap\,\{\,N\smallsetminus M:(M,N)\in\omega\,\}\cap V(T^{\prime}).$
@@
Suppose first that $\omega$ is empty. Then $V(T^{\prime})$ is empty. Hence $T$ is a $K_{2}$ both whose endvertices have triangle or 4-cycle torsos that represent links. By Lemma˜9.4, both torsos are triangles. Hence the corner $AC$ is potter, so $A\cap C=2$, contradicting that $E\subseteq A\cap C$ where $E$ has size $\geqslant 4$.
@@
To show that $(E,F)\leqslant$ the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$, by Lemma˜9.46 and Lemma˜5.7 it remains to show $E\subseteq V(G_{t^{*}})$. Assume for a contradiction that there is a vertex $v\in E\smallsetminus V(G_{t^{*}})$. Recall that $E\subseteq A\cap C$, so in particular $E\subseteq V(\beta)$. Hence there is a node $s$ of $T$ with $v\in V_{s}$, and $s\neq t^{*}$ by assumption. Let $r$ be the penultimate node on the path $sTt^{*}$.
[…]
@@
Assume for a contradiction that $\sigma$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$. By Lemma˜9.48, there is a 3-connected torso $H_{t}$ of $\mathcal{O}$ such that $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, say, where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ denotes the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
Assume first that $H_{t}$ is good. Then $(U_{t},W_{t})$ is a totally-nested tetra-separation by ˜9.41. In particular, $(U_{t},W_{t})\in\sigma$. Since $(E,F)$ interlaces $\sigma$, we have one of $(U_{t},W_{t})<(E,F)$ and $(U_{t},W_{t})<(F,E)$. We cannot have $(U_{t},W_{t})<(F,E)$, since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t})$ contradicts the fact that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by ˜5.2. Hence $(U_{t},W_{t})<(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$. Since $(U_{t},W_{t})$ is the left-reduction of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
Assume otherwise that $G_{t}$ is bad. Hence $G_{t}$ has four or five vertices. If $H_{t}$ has four vertices, then all but at most one of the four vertices in $S(U^{\prime}_{t},W^{\prime}_{t})$ has $\geqslant 2$ neighbours in $W^{\prime}_{t}\smallsetminus U^{\prime}_{t}$. If $H_{t}$ has five vertices, then all four vertices in $S(U^{\prime}_{t},W^{\prime}_{t})$ have $\geqslant 2$ neighbours in $W^{\prime}_{t}\smallsetminus U^{\prime}_{t}$. In either case, we have $|U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}|\leqslant 1$. Since $(E,F)\leqslant(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, we have $E\smallsetminus F\subseteq U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$. Therefore, $|E\smallsetminus F|\leqslant 1$, which by Lemma˜5.5 contradicts that $(E,F)$ is a tetra-separation. ∎
@@
By the (Crossing Lemma). (4.1), all links have size two. By Lemma˜9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$. By ˜9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. By Lemma˜9.49, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$. Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$. Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$. The torso $\tau$ is 4-connected by Lemma˜8.36. We can obtain $\tau$ from $\mathcal{O}$ by replacing each good torso with a $K_{4}$, leaving only 3-connected torsos with five or four vertices, and contracting some free edges in torsos that are 4-cycles or triangles. Each torso that is a 4-cycle neighbours two 3-connected torsos by Lemma˜9.4, and the 4-cycle-torso witnesses that both 3-connected torsos are good. Hence the free edges of the 4-cycle-torsos are contained in the separators of the elements of $\sigma(\mathcal{O})$ that stem from the good neighbouring torsos, and so they are contracted. Hence $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. ∎